40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 14 03:42:31 2015, Jenne, Update, LSC, Thoughts on going forward with variable finesse 
    Reply  Thu Jan 15 04:41:01 2015, Jenne, Update, LSC, Thoughts on going forward with variable finesse 
       Reply  Thu Jan 15 14:28:14 2015, Jenne, Update, LSC, LSC model change idea SwitchableErrorSignalProposal_Jan2015.pdf
          Reply  Fri Jan 16 03:31:35 2015, Jenne, Update, LSC, LSC model change implemented LSCorig_15Jan2015.pngLSCpost_15Jan2015.png
             Reply  Fri Jan 16 18:46:15 2015, Koji, Update, LSC, LSC model change implemented 
                Reply  Fri Jan 16 20:01:32 2015, Jenne, Update, LSC, LSC model change implemented 
             Reply  Tue Jan 20 15:09:01 2015, Jenne, Update, LSC, LSC model change implemented 
             Reply  Sat Mar 7 18:53:48 2015, Jenne, Update, LSC, DoF selector matrices replaced with filter banks DoFselectors_7March2015.png
Message ID: 10923     Entry time: Tue Jan 20 15:09:01 2015     In reply to: 10910
Author: Jenne 
Type: Update 
Category: LSC 
Subject: LSC model change implemented 

 

Quote:

Brain not working anymore now that it's ~4am, but I need to rethink and recheck to make sure that the PD whitening triggering is still okay and working.  Or maybe we can remove it, and just include that in the scripts, as Rana has been suggesting for ages.  Thoughts for tomorrow.

LSC whitening triggering was not working, because of the implementation of the double-rows for the input matrix.  I have modified the c-code that looks at the input matrix and triggers, and decides when to turn on the PD whitening, so that it now works.

ELOG V3.1.3-